OpenStudy (anonymous):

dont get it 5. Write the name of each property illustrated by the expressions below: a. 823 x 1 = 823 b. 96 + 60 = 12(8 + 5) c. 3 x (4 x 5) = (3 x 4) x 5 d. 3 + (4 + 5) = 3 + (5 + 4) e. 0 + 11 = 11 f. 96(75) = 75(96)

OpenStudy (anonymous):

a. 823 x 1 = 823 identity property on multiplication

OpenStudy (anonymous):

so like the answer ?

OpenStudy (anonymous):

e. 0 + 11 = 11 identity property on addition

OpenStudy (anonymous):

oooo

OpenStudy (anonymous):

c. 3 x (4 x 5) = (3 x 4) x 5 associative property

OpenStudy (anonymous):

lol was gonna say that one

OpenStudy (anonymous):

on multiplication

OpenStudy (anonymous):

3 + (4 + 5) = 3 + (5 + 4) associative property on addition

OpenStudy (anonymous):

b. 96 + 60 = 12(8 + 5) commutative property

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

96 + 60 = 12(8 + 5) , i think this is distributive property over addition

OpenStudy (anonymous):

96(75) = 75(96) commutative over multiplication

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw